Find the area under the standard normal curve to the right of -1.24

Group of answer choices

A. 0.124

B. 0.89

C. 0.1

D. 1.24
Answer #1

Solution:

Using standard normal table,

P(z > -1.24) = 1 - P(z < -1.24) = 1 - 0.1075 = 0.89

Area = 0.89